The time it takes for a motorist to press the brakes after seeing a red light is uniformly distributed between 1 and 3 seconds. Find the probability that it takes less than 2 seconds for a randomly selected motorist to press the brakes after seeing a red light.
Answer #1

Solution :

Given that,

a = 1

b = 3

P(x < c) = (c - a) / (b - a)

P(x < 2) = (2 - 1) / (3- 1) =0.5

probability =0.5